OpenStudy (anonymous):

If f(x) = 5x - 9, find the value of f(-2). A. 8 B. -19 C. 23 D. -23

OpenStudy (anonymous):

B

OpenStudy (anonymous):

f(-2) = 5(-2) -9 = -10 -9 = -19

OpenStudy (anonymous):

put x=-2 in the equation you will get it

OpenStudy (anonymous):

ok